# OptiMind: Teaching LLMs to Think Like Optimization Experts

### v1

@@ Abstract
Mathematical programming – the task of expressing operations and decision-making problems in precise mathematical language – is fundamental across domains, yet remains a skill-intensive process requiring operations research expertise. Recent advances in large language models for complex reasoning have spurred interest in automating this task, translating natural language into executable optimization models. Current approaches, however, achieve limited accuracy, hindered by scarce and noisy training data without leveraging domain knowledge. In this work, we systematically integrate optimization expertise to improve formulation accuracy for mixed-integer linear programming, a key family of mathematical programs. Our approach first cleans training data through class-based error analysis to explicitly prevent common mistakes within each optimization class. We then develop multi-turn inference strategies that guide LLMs with class-specific error summaries and solver feedback, enabling iterative refinement. Experiments across multiple base LLMs demonstrate that combining cleaned data with domain-informed prompting and feedback improves formulation accuracy by 14 percentage points on average, enabling further progress toward robust LLM-assisted optimization formulation.
@@ 1 Introduction
Researchers and practitioners have begun exploring whether large language models (LLMs) can automate the formulation task: translating natural language problem descriptions into executable optimization models (Ramamonjison et al., 2022; Tang et al., 2024; Yang et al., 2025). Success on this front would lower a major barrier to broader use of optimization, democratizing its benefits. Yet current systems remain far from this goal. Accuracy is limited by training data quality, including lack of diversity and frequent errors from synthetic data generation. On the inference side, most existing approaches rely on a simple prompt that includes the natural language problem description and the modeling goal (e.g., generate an integer programming formulation in Pyomo, OR-Tools, or GurobiPy). This leaves out more structured prompting strategies that may yield stronger outcomes. The problem is further compounded by noisy benchmarks, such as ambiguous questions, missing data, and incorrect ground-truth values, where insufficient manual cleaning produces high error rates that obscure true performance. Critically, existing approaches underutilize domain expertise in benchmark quality control, training, and inference.
@@
In this work, we study how optimization expertise can be systematically integrated – both at training time and inference time – to improve formulation accuracy at scale. Our first step is to clean existing training data. Motivated by our observation that existing LLMs often repeat similar mistakes within each optimization problem class, we first map problems into a fixed set of canonical classes (e.g., set cover, flow shop scheduling), then manually analyze a small subset of representative examples from each class to extract common formulation errors. We then use this error analysis to clean the entire training dataset by explicitly prompting strong LLMs to follow class-specific hints and avoid common error patterns, further using self-consistency to improve solution generation quality.
Beyond training data, we posit that achieving high accuracy in optimization formulation requires advanced inference techniques. Our error analyses and formulation hints derived from the training data are effectively incorporated into prompts to improve generation quality. Based on our observation that LLMs often generate noisy outputs, producing infeasible models or code with execution errors, we extend the inference pipeline into a multi-step process. In this pipeline, solver feedback – such as infeasibility detection, execution errors, and the previous optimal solution – is incorporated in conjunction with the hints to iteratively refine outputs, with self-consistency applied at each step to further stabilize performance. We henceforth refer to our training and inference framework as OptiMind, as it leverages human reasoning to improve a critical optimization task.
@@
We evaluate our approach on the recently released gpt-oss-20b, applying both training-data and inference enhancements. To ensure reliable evaluation, we carefully re-clean three of the most challenging public benchmarks. On average, OptiMind improves performance by 14 percentage points across the evaluated benchmarks compared to the base model. An ablation study confirms that both training-data and inference components contribute significantly to these improvements. We further show that incorporating hints improves accuracy across multiple models, pointing to the robustness of our approach. Our contributions are summarized as follows:
@@
  • We propose a multi-turn inference pipeline that integrates class-specific error summaries with self-consistency and solver feedback to iteratively refine LLM outputs.
  • We conduct extensive empirical evaluation on three manually cleaned benchmarks, showing that our OptiMind framework – combining enhanced training data, domain expertise, and feedback-guided inference – substantially improves formulation accuracy achieving state-of-the-art results.
@@ 3 Related Work
A growing body of work focuses on building benchmarks for natural language to optimization formulation translation. NL4LP (Ramamonjison et al., 2022) was one of the early benchmarks introduced in a NL4OPT competition, focusing on Linear Programming (LP) problems. Mamo (Huang et al., 2024) expands the scope to MILPs, with Mamo Easy and Mamo Complex reflecting different difficulty levels. NLP4LP (AhmadiTeshnizi et al., 2024) contains LP and ILP problems collected from textbooks and lecture notes. Xiao et al. (2024) released ComplexOR, a benchmark of OR problems collected from both industrial and academic scenarios. IndustryOR (Tang et al., 2024) provides 100 real-world OR problems across eight industries, while OptiBench (Yang et al., 2025) extends coverage to nonlinear and tabular problems. OptMATH (Lu et al., 2025) further introduces GPT-synthesized benchmark with longer natural language contexts and complex constraints. A major limitation in current benchmarks is their high error rates. Both our own experience and a recent survey (Xiao et al., 2025) reveal frequent issues, with up to 30–50% error rates, including missing data, ambiguous formulations, and incorrect ground-truth answers. We thus have dedicated significant efforts to cleaning some of the benchmarks.
Several prompting strategies have been developed to improve formulation accuracy. Chain-of-experts (Xiao et al., 2024) and OptiMUS (AhmadiTeshnizi et al., 2024) adopt agentic frameworks that split the task into specialized LLM agents for modeling, programming, and evaluation. Search-based methods such as AutoFormulation (Astorga et al., 2025) use Monte-Carlo Tree Search to separately construct variables, constraints, and objectives, while pruning equivalent formulations at the symbolic level. More recently, OptiTree (Liu et al., 2025) further decomposes complex optimization tasks into simpler subproblems to improve the overall solution. Multiple works have explored fine-tuning strategies for optimization datasets. ORLM (Tang et al., 2024), ReSocratic (Yang et al., 2025), and OptMATH (Lu et al., 2025) apply supervised fine-tuning (SFT) on LLM-synthesized datasets to improve performance on IndustryOR, OptiBench, and OptMATH, respectively, while LLMOpt (Jiang et al., 2025) uses Kahneman-Tversky Optimization (KTO) on synthetic data. SIRL (Chen et al., 2025) combines SFT with RL to further boost results. However, reproducibility of these works remains a challenge: some models are highly prompt-sensitive and complete training data and latest checkpoints are often unavailable.
@@ 4 Method
4.1 Test Data Cleaning
We consider three of the most challenging optimization formulation benchmarks – IndustryOR, Mamo Complex, and OptMATH – where even the strongest models at the time only report accuracies up to 20%–50% (Jiang et al., 2025; Liu et al., 2025; Lu et al., 2025). By contrast, earlier LP-centric benchmarks (e.g., NL4LP (Ramamonjison et al., 2022), NLP4LP (AhmadiTeshnizi et al., 2024)) already see 90–95%+ accuracy (see e.g. (Liu et al., 2025; Lu et al., 2025)), highlighting substantial headroom on these harder tasks. This raises a natural question: where do existing models fall short, and how can we improve?
@@ 4.2 Problem-Class Specific Error Analysis
While highly capable LLMs perform strongly on the cleaned benchmarks, in this work we take the ambitious goal of further improving their performance in optimization formulation. Upon examining their outputs on a small training subset, we find that formulation mistakes still occur, and often similar types of mistakes recur within the same optimization category: for example, in the Traveling Salesman Problem (TSP), the LLMs often mess up the subtour elimination constraint by incorrectly applying it to the fixed starting node (Fig. 4, middle box). Similarly, we see frequent mistakes in the signs of flow-conservation constraints in network-flow and inventory management problems; for example, if $B[u]$ denotes the net supply of a node $u$ (i.e., positive if there is extra supply), the flow conservation equation should read [outflow from $u$] - [inflow into $u$] =B[u]=B[u], but we often found the outflow and inflow terms swapped. This suggests a simple yet intuitive intervention: summarize short, targeted hints that capture the most common failure modes in each optimization category and attach the appropriate hint when solving problems from that category.
To this end, we develop a class-based error analysis. We classify all problems in the open-source optimization training sets (OptMATH (Lu et al., 2025) and OR-Instruct (Tang et al., 2024), which is the training data corresponding to the IndustryOR benchmark) into the 50 seed classes defined in the OptMATH problem generators; these classes provide a wide coverage of typical MILP-type problems. For each class, we run gpt-oss-20b to produce answers and select 10–20 instances where its answer disagrees with the original ground-truth label. This label mismatch can be attributed either to model errors in the solution generation or quality issues with the questions themselves (details in §4.3). Our optimization experts manually review each item to identify the source of the disagreement; when the model is at fault, we write a short error summary and a concrete hint that would have prevented the mistake (see Fig. 4 (left) for an example). We then aggregate these annotations into a dictionary that maps each problem type into a list of (error summary, preventive hint) pairs.
@@ 4.3 Training Data Cleaning
  • Solution regeneration using error analysis and majority voting. Here we employ a simplified version of our inference process described in more detail in §4.5. That is, we use gpt-oss-20b augmented with the class-specific error summaries and hints described in §4.2 to reduce recurrent modeling mistakes. We use majority vote with K=64K{=}64 samples, yielding higher-quality solutions.
@@
  • Resolve ambiguity with expert edits. We manually clean classes with systematic ambiguity (e.g., Job Shop); an example is in Appendix A.8. In general, such ambiguity cannot be resolved reliably in a fully automated way; broader coverage of additional types is left to future work.
@@ 4.4 Supervised Fine-Tuning
We use supervised fine-tuning (SFT) to strengthen the model’s formulation and coding ability. Concretely, we fully fine-tune gpt-oss-20b on our cleaned training dataset. Let the SFT dataset be $\mathcal{D}_{\mathrm{SFT}}={(x_{i},y_{i})}_{i=1}^{N}$, where $x_{i}$ is the problem description and $y_{i}$ is the completion sequence formed by concatenating the model’s thinking tokens, the mathematical formulation, and the solver code. We train with standard sequence-to-sequence loss: $\mathcal{L}_{\mathrm{SFT}}(\theta)=-\mathbb{E}_{(x,y)\sim\mathcal{D}_{\mathrm{SFT}}}\sum_{t=1}^{|y|}\log p_{\theta}\bigl(y_{t}\mid y_{<t},x\bigr)$, where $\theta$ denotes all trainable parameters of the model; $p_{\theta}(\cdot)$ is the model’s output distribution; $y_{t}$ is the target token at position $t$; and $y_{<t}:=(y_{1},\ldots,y_{t-1})$ are the prefix of previously generated target tokens. Full training details are provided in the Appendix A.6.1.
@@ 4.5 Inference
We adopt a three-stage inference pipeline with (i) error-aware prompting, (ii) self-consistency via majority voting, and (iii) multi-turn correction with tool feedback. As shown in §5, these components reinforce each other to form a robust pipeline for solving optimization problems.
Improved prompts with error-analysis-based hints. We incorporate the class-based error analysis from §4.2 into the inference process. The model first classifies each test instance into one of the 50 classes that were defined in the training data error analysis, then augments its prompt with all error–hint pairs from that class to guide solution generation. Since the training instances from the same class can differ in their characteristics (e.g., some multi-period inventory problems allow backorders while others do not), we clarify in the prompt that hints should be applied only when relevant, allowing the model to ignore those hints that do not fit the problem description. This category-based targeted application helps the model avoid recurrent pitfalls and reduces common mistakes. On top of the hints we obtained from our analysis, we add general guidelines for correct formulation, which are derived from our optimization experts’ experience.
Self-consistency with majority voting. We sample multiple reasoning traces and take the self-consistent answer that appears most frequently. Taking the majority voting solution dampens stochastic variance from sampling, stabilizes final predictions, and yields strong gains even under otherwise simple prompts.
Multi-turn correction. We then run a multi-turn loop to further reduce errors and leverage the model’s self-improvement. After each round, we extract and execute the generated Python code, and collect system feedback containing Gurobi logs (e.g., the optimal solution found or whether infeasibility was detected) or execution errors (when code fails to run). We feed this feedback to the model, ask it to assess correctness, and, if needed, revise the code. With strong reasoning models such as gpt-oss, most low-level coding errors are fixed within a few turns, and the model can also self-correct modeling issues. Appendix A.5 provides an example of nontrivial self-corrections.
Putting it together, the full inference pipeline proceeds as follows: the initial turn classifies the problem; the next turn solves it using the returned hints for the predicted class(es). We generate $K$ solutions and select the majority-vote answer. We then return the solutions and system feedback to the model for self-correction, again aggregating by majority across multiple samples per turn. This correction loop is repeated for $M$ rounds. The pseudocode is provided in Alg. 1 in Appendix A.4.

### v2

@@ Abstract
Mathematical programming – the task of expressing operations and decision-making problems in precise mathematical language – is fundamental across domains, yet remains a skill-intensive process requiring operations research expertise. Recent advances in large language models for complex reasoning have spurred interest in automating this task, translating natural language into executable optimization models. Current approaches, however, achieve limited accuracy, hindered by scarce and noisy training data without leveraging domain knowledge. In this work, we systematically integrate optimization expertise to improve formulation accuracy for mixed-integer linear programming, a key family of mathematical programs. Our OptiMind framework leverages semi-automated, class-based error analysis to guide both training and inference, explicitly preventing common mistakes within each optimization class. Our resulting fine-tuned LLM significantly improves formulation accuracy by 20.7% across multiple optimization benchmarks, with consistent gains under test-time scaling methods such as self-consistency and multi-turn feedback, enabling further progress toward robust LLM-assisted optimization formulation.
@@ 1 Introduction
Researchers and practitioners have begun exploring whether large language models (LLMs) can automate the formulation task: translating natural language problem descriptions into executable optimization models (Ramamonjison et al., 2022; Tang et al., 2024; Yang et al., 2025b). Success on this front would lower a major barrier to broader use of optimization, democratizing its benefits. Yet current systems remain far from this goal. Accuracy is limited by training data quality, including lack of diversity and frequent errors from synthetic data generation. On the inference side, most existing approaches rely on a simple prompt that includes the natural language problem description and the modeling goal (e.g., generate an integer programming formulation in Pyomo, OR-Tools, or GurobiPy). This leaves out more structured prompting strategies that may yield stronger outcomes. The problem is further compounded by noisy benchmarks, such as ambiguous questions, missing data, and incorrect ground-truth values, where insufficient manual cleaning produces high error rates that obscure true performance. Critically, existing approaches underutilize domain expertise in benchmark quality control, training, and inference.
@@
In this work, we study how optimization expertise can be systematically integrated – both at training time and inference time – to improve formulation accuracy at scale. A core component of achieving strong training outcomes is high-quality data. Motivated by our observation that existing LLMs often repeat similar mistakes within each optimization problem class, we first map problems into a fixed set of canonical classes (e.g., set cover, flow shop scheduling), then manually analyze a small subset of representative examples from each class to extract common formulation errors. We then leverages this error analysis to automate the training dataset cleaning pipeline by explicitly prompting strong LLMs to follow class-specific hints and avoid common error patterns, further using self-consistency to improve solution generation quality.
These class-based error analyses not only guide the fine-tuning of strong LLMs, but they can also improve performance at inference: the same formulation hints derived from the training data can be incorporated directly into prompts and help reduce common formulation errors. As LLMs may still generate noisy outputs, producing infeasible models or code with execution errors, we further apply test-time scaling methods such as self-consistency and multi-step refinement when additional compute is available. We henceforth refer to our combined training and inference framework as OptiMind, as it leverages domain expertise to improve a critical optimization task.
@@
We evaluate our approach on the recently released gpt-oss-20b, applying both training-data and inference enhancements. To ensure reliable evaluation, we carefully re-clean three of the most challenging public benchmarks. Across these benchmarks, OptiMind improves absolute accuracy over the gpt-oss-20b base model by between $13\%$ and $21\%$ and consistently outperforms other open-source models of comparable or larger size, while approaching the performance of proprietary frontier models. An ablation study confirms that both training-data and inference components contribute significantly to these improvements. We further show that incorporating hints improves accuracy across multiple models, pointing to the robustness of our approach. Our contributions are summarized as follows:
@@
  • We train a 20B-parameter gpt-oss-20b variant on the semi-automatically cleaned data, yielding a strong open-source model for optimization formulation.
  • We propose a inference pipeline that integrates class-specific error summaries as error-analysis-based hints, enabling optional test-time scaling with majority voting and self-correction.
  • We conduct extensive empirical evaluation on three manually cleaned benchmarks, showing that our OptiMind framework substantially improves formulation accuracy, surpasses all open-source baselines of similar or larger size, and achieves performance competitive with much larger proprietary frontier models.
@@ 3 Related Work
A growing body of work focuses on building benchmarks for natural language to optimization formulation translation. NL4LP (Ramamonjison et al., 2022) was one of the early benchmarks introduced in a NL4OPT competition, focusing on Linear Programming (LP) problems. Mamo (Huang et al., 2024) expands the scope to MILPs, with Mamo Easy and Mamo Complex reflecting different difficulty levels. NLP4LP (AhmadiTeshnizi et al., 2024) contains LP and ILP problems collected from textbooks and lecture notes. Xiao et al. (2024) released ComplexOR, a benchmark of OR problems collected from both industrial and academic scenarios. IndustryOR (Tang et al., 2024) provides 100 real-world OR problems across eight industries, while OptiBench (Yang et al., 2025b) extends coverage to nonlinear and tabular problems. OptMATH (Lu et al., 2025) further introduces GPT-synthesized benchmark with longer natural language contexts and complex constraints. LogiOR (Yang et al., 2025a) proposes a new optimization modeling benchmark from the logistics domain, containing more complex problems with standardized annotations.
A major limitation in current benchmarks is their high error rate. Both our own experience and a recent survey (Xiao et al., 2025b) reveal frequent issues, including missing data, ambiguous formulations, and incorrect ground-truth answers. We thus have dedicated significant efforts to cleaning some of the benchmarks. In parallel to our efforts, several recent works also focus on improving the quality of existing datasets. SIRL (Chen et al., 2025) releases cleaned versions of NL4OPT, IndustryOR, MAMO-ComplexLP, and MAMO-EasyLP. Yang et al. (2025a) further provide cleaned variants of IndustryOR, ComplexOR, and NL4LP as part of the LogiOR benchmark. Xiao et al. (2025a) examine six mainstream benchmarks, including IndustryOR and MAMO-Complex, and find that about 30-50% of the data are problematic; they discard these instances and release a curated subset containing only validated problems.
Several prompting strategies have been developed to improve formulation accuracy. Chain-of-experts (Xiao et al., 2024) and OptiMUS (AhmadiTeshnizi et al., 2024) adopt agentic frameworks that split the task into specialized LLM agents for modeling, programming, and evaluation. Search-based methods such as AutoFormulation (Astorga et al., 2025) use Monte-Carlo Tree Search to separately construct variables, constraints, and objectives, while pruning equivalent formulations at the symbolic level. More recently, OptiTree (Liu et al., 2025) further decomposes complex optimization tasks into simpler subproblems to improve the overall solution. Multiple works have explored fine-tuning strategies for optimization datasets. ORLM (Tang et al., 2024), ReSocratic (Yang et al., 2025b), and OptMATH (Lu et al., 2025) apply supervised fine-tuning (SFT) on LLM-synthesized datasets to improve performance on IndustryOR, OptiBench, and OptMATH, respectively, while LLMOpt (Jiang et al., 2025) uses Kahneman-Tversky Optimization (KTO) on synthetic data. SIRL (Chen et al., 2025) combines SFT with RL to further boost results. However, reproducibility of these works remains a challenge: some models are highly prompt-sensitive and complete training data and latest checkpoints are often unavailable.
@@ 4.1 Problem-Class Specific Error Analysis
While highly capable LLMs already show strong potential for optimization formulation, in this work we take on the ambitious goal of further improving their performance. Upon examining their outputs on a small training subset, we find that formulation mistakes still occur, and often similar types of mistakes recur within the same optimization category: for example, in the Traveling Salesman Problem (TSP), the LLMs often mess up the subtour elimination constraint by incorrectly applying it to the fixed starting node (Fig. 4, middle box). Similarly, we see frequent mistakes in the signs of flow-conservation constraints in network-flow and inventory management problems; for example, if $B[u]$ denotes the net supply of a node $u$ (i.e., positive if there is extra supply), the flow conservation equation should read [outflow from $u$] - [inflow into $u$] =B[u]=B[u], but we often found the outflow and inflow terms swapped. This suggests a simple yet intuitive intervention: summarize short, targeted hints that capture the most common failure modes in each optimization category and attach the appropriate hint when solving problems from that category.
To this end, we develop a class-based error analysis. We classify all problems in the open-source optimization training sets (OptMATH (Lu et al., 2025) and OR-Instruct (Tang et al., 2024), which is the training data corresponding to the IndustryOR benchmark) into the 53 seed classes defined in the OptMATH problem generators; these classes provide a wide coverage of typical MILP-type problems. For each class, we run gpt-oss-20b to produce answers and select 10–20 instances where its answer disagrees with the original ground-truth label. This label mismatch can be attributed either to model errors in the solution generation or quality issues with the questions themselves (details in §4.2). Our optimization experts manually review each item to identify the source of the disagreement; when the model is at fault, we write a short error summary and a concrete hint that would have prevented the mistake (see Fig. 4 (left) for an example). We then aggregate these annotations into a dictionary that maps each problem type into a list of (error summary, preventive hint) pairs.
@@ 4.2 Training Data Cleaning
  • Solution regeneration using error analysis and majority voting. Here we employ a simplified version of our inference process described in more detail in §4.4. That is, we use gpt-oss-120b augmented with the class-specific error summaries and hints described in §4.1 to reduce recurrent modeling mistakes. We use majority vote with K=64K{=}64 samples, yielding higher-quality solutions.
@@
  • Regenerate and clarify OptMATH instances. Following the data generation process of OptMATH (Lu et al., 2025), we re-generate 53 instances for each seed class and back-translate the problem descriptions using the gpt-oss-120b model. Our optimization experts manually inspect 159 regenerated instances (three per class) to confirm that key parameters are present and consistent. Training on this regenerated data improved downstream performance compared to using the original OptMATH problems; see our ablation results in §5.3.
@@ 4.3 Supervised Fine-Tuning
We use supervised fine-tuning (SFT) to strengthen the model’s formulation and coding ability. Concretely, we fully fine-tune gpt-oss-20b on our cleaned training dataset. Let the SFT dataset be $\mathcal{D}_{\mathrm{SFT}}={(x_{i},y_{i})}_{i=1}^{N}$, where $x_{i}$ is the problem description and $y_{i}$ is the completion sequence formed by concatenating the model’s thinking tokens, the mathematical formulation, and the solver code. We train with standard sequence-to-sequence loss: $\mathcal{L}_{\mathrm{SFT}}(\theta)=-\mathbb{E}_{(x,y)\sim\mathcal{D}_{\mathrm{SFT}}}\sum_{t=1}^{|y|}\log p_{\theta}\bigl(y_{t}\mid y_{<t},x\bigr)$, where $\theta$ denotes all trainable parameters of the model; $p_{\theta}(\cdot)$ is the model’s output distribution; $y_{t}$ is the target token at position $t$; and $y_{<t}:=(y_{1},\ldots,y_{t-1})$ are the prefix of previously generated target tokens. Full training details are provided in the Appendix A.7.
@@ 4.4 Inference
We adopt an inference pipeline with error-aware prompting as the default, and, when additional computation budget is available, we optionally apply two test-time scaling techniques: (1) self-consistency via majority voting and (2) multi-turn correction with tool feedback. As shown in §5, these components reinforce each other to form a robust pipeline for solving optimization problems.
Improved prompts with error-analysis-based hints. We incorporate the class-based error analysis from §4.1 into the inference process. The model first classifies each test instance into one of the 53 classes that were defined in the training data error analysis, then augments its prompt with all error–hint pairs from that class to guide solution generation. Since the training instances from the same class can differ in their characteristics (e.g., some multi-period inventory problems allow backorders while others do not), we clarify in the prompt that hints should be applied only when relevant, allowing the model to ignore those hints that do not fit the problem description. This category-based targeted application helps the model avoid recurrent pitfalls and reduces common mistakes. On top of the hints we obtained from our analysis, we add general guidelines for correct formulation, which are derived from our optimization experts’ experience.
While the hinting mechanism is our main contribution on the inference side, we also study two standard test-time scaling techniques that can further improve performance when additional computation is available.
Self-consistency with majority voting. Self-consistency with majority voting samples multiple reasoning traces for the same instance and returns the answer that appears most frequently, reducing sampling variance and providing modest accuracy gains at the cost of extra samples.
Multi-turn correction. Multi-turn correction runs a short feedback loop in which we execute the generated Python code, collect solver logs or execution errors, feed this feedback back to the model, and let it revise its formulation; with a few turns, most coding errors and some modeling issues are corrected, but each additional turn increases inference cost. Appendix A.6 provides an example of nontrivial self-corrections.
Putting it together, the full inference pipeline proceeds as follows: the initial turn classifies the problem; the next turn solves it using the returned hints for the predicted class(es); we may then generate $K$ solutions and select the majority-vote answer, and, if further compute budget is available, repeat this correction loop for $M$ rounds using tool feedback, as summarized in Alg. 1 in Appendix A.5.
4.5 Ensuring Reliable Evaluation via Careful Test Data Cleaning
We consider three of the most challenging optimization formulation benchmarks – IndustryOR, Mamo Complex, and OptMATH – where even the strongest models at the time only report accuracies up to 20%–50% (Jiang et al., 2025; Liu et al., 2025; Lu et al., 2025). By contrast, earlier LP-centric benchmarks (e.g., NL4LP (Ramamonjison et al., 2022), NLP4LP (AhmadiTeshnizi et al., 2024)) already see 90–95%+ accuracy (see e.g. (Liu et al., 2025; Lu et al., 2025)), highlighting substantial headroom on these harder tasks.